OpenStudy (pottersheep):

Grade 11 Math help PLEASE!!

OpenStudy (pottersheep):

The sum of the first 5 terms in an arithmetic series is 0. The last term is 4. Determine the first term and the common difference. The answer is supposed to be 4, -1.

OpenStudy (pottersheep):

I don't know how to find a & d or anything :(

OpenStudy (anonymous):

you will have to apply some formula Sum formula \[S_n=n/2 (a_1+a_2)\]

OpenStudy (anonymous):

*correction \[S_n=n/2 (a_1+a_n)\]

OpenStudy (anonymous):

so what's given?

OpenStudy (pottersheep):

I have

OpenStudy (pottersheep):

5 terms, sum of 5 terms = 0 , term 5 = 4

OpenStudy (anonymous):

good! so n=5 \[S_n=0\] \[a_n=4\]

OpenStudy (pottersheep):

The formula we use is Sn = n/2 (2a + (n-1)d)

OpenStudy (pottersheep):

hmm..

OpenStudy (anonymous):

I think you are better off using my formula \[S_n=n/2 (a_1+a_n)\] plug in numbers \[0= 5/2 (a_1 +4)\] solve for \[a_1\]

OpenStudy (pottersheep):

I'll try it out , thank you ~=)